OpenStudy (anonymous):

You roll two dice, what is the probability that the sum of the dice is less than 5

OpenStudy (mathmale):

Why not write out all of the possibilities? First die: Get a ' 1 '; on the second die a 1, 2, 3, 4, 5, or 6. How many sums of ' 1 ' and the other number is less than 5? First die: Get a '2 '; and so on. count the number of possibilities for which the sum would be less than 5. this count, divided by the number of ALL possible outcomes, will be the probability you seek.

OpenStudy (anonymous):

there are two possibilities of getting a 2 and 1dice 1 =1 dice 2=2 OR dice 1=1 dice 2=2 there is 36 different possibilities so 2/36 is probability simplify 2/36

OpenStudy (mathmale):

I'd still advise writing out all the possible outcomes, and then counting how many of the sums are less than 5 (that is, are 2, 3 or 4. A 1 on one die and a 1 on the 2nd would count: sum is 2. A 1 on one die and a 2 on the next would count: sum is 3. A 1 on one die and a 3 on the next would count: sum is 4. STOP. Now: A 2 on one die and a 1 on the 2nd would count: sum is 3. A 2 on one die and a 2 on the 2nd would also count: sum is 4. Stop A 3 on one die .... and so on. @beautifulmelodies: are you sure 2/36 is correct? Discuss how you got this result.

OpenStudy (anonymous):

@mathmale There is this thing at school, called math class. You go there and you learn. Different teachers' teacher different way's. :)

OpenStudy (mathmale):

@jamjam11: We need to hear from YOU. What have you done so far? Pleare it with us, so that we could give you appropriate feedback.

OpenStudy (mathmale):

@beautifulmelodies: Doesn't matter how your learned the material. I'm asking you to explain how you got your solution and to justify it.

OpenStudy (anonymous):

Why? Because your mad I am helping someone? I have been wathcing you latley. You don't seme like a good sport. I don't have to explain anything to you. If JamJam11 want's to know how I go the answer. Then I would be happy to explain it to HIM.

OpenStudy (the_fizicx99):

Lol, wtf..

OpenStudy (anonymous):

Please @mathmale let's be ploite to one another, and not argue on question's :)

OpenStudy (the_fizicx99):

Um you copied that off yahoo. Just saying'.

OpenStudy (anonymous):

No, I didn't. I learned possibilities. Just today at school. I am not a cheater.

OpenStudy (the_fizicx99):

Mhmm, k.

OpenStudy (anonymous):

I have you know, that spam is against the OpenStudy rules. Unless your going to help JamJam11 I suggest you stop.

OpenStudy (mathmale):

There are 6 possible outcomes for the throw of the first die, and 6 for the throw of the second die. Therefore, as B. M. says, there are 36 possible outcomes. However, we're interested ONLY in those that add up to 2, 3 or 4. All others must be rejected. 1+1 = 2 OK 1+2 = 3 OK 1+3 = 4 OK 2+1 is really the same as 1 + 2, but let's count it anyway, at least for now. 2+1 = 3 OK (but this duplicates 1+2, doesn't it?) 2+2 = 4 OK 3+1 = ? So I could at least 6 unique outcomes whose sums are less than 5. So that automatically rules out an answer of 2/36. I'd go for 6/36, or 1/6, myself, but am open to arguments by anyone who thinks the correct result is different.

OpenStudy (the_fizicx99):

Well that was a pitiful attack, it's not spam I only commented 3 times, related to the subject.

OpenStudy (mathmale):

Friends, let's concentrate on the Math alone for now, OK? Thanks.

OpenStudy (anonymous):

The possible ways the 2 dice will roll to 5 are 1,3; 2,2; and ,3,1. So there is 3 way's So the probability is 3 in 36 or 1/12 I just double checked my work. My bad.